# Succeeding at Scale: Automated Multi-Retriever Fusion and Query-Side Adaptation for Multi-Tenant Search

### v1

@@ Abstract
Large-scale multi-tenant retrieval systems amass vast user query logs yet critically lack the curated relevance labels required for effective domain adaptation. This "dark data" problem is exacerbated by the operational cost of model updates: jointly fine-tuning query and document encoders requires re-indexing the entire corpus, which is prohibitive in multi-tenant environments with thousands of isolated indices. To address these dual challenges, we introduce DevRev Search, a passage retrieval benchmark for technical customer support constructed through a fully automatic pipeline. We employ a fusion-based candidate generation strategy, pooling results from diverse sparse and dense retrievers, and utilize an LLM-as-a-Judge to perform rigorous consistency filtering and relevance assignment. We further propose a practical Index-Preserving Adaptation strategy: by fine-tuning only the query encoder via Low-Rank Adaptation (LoRA), we achieve competitive performance improvements while keeping the document index frozen. Our experiments on DevRev Search and SciFact demonstrate that targeting specific transformer layers in the query encoder yields optimal quality-efficiency trade-offs, offering a scalable path for personalized enterprise search.
Succeeding at Scale: Automated Multi-Retriever Fusion and Query-Side Adaptation for Multi-Tenant Search
Prateek Jain1thanks: Corresponding author: prateek.jain@devrev.ai, Shabari S Nair2, Ritesh Goru1, Prakhar Agarwal3,
Ajay Yadav2, Yoga Sri Varshan Varadharajan2, Constantine Caramanis2 1DevRev, Austin, USA 3DevRev, Bengaluru, India 2The University of Texas at Austin, Austin, USA
@@ 1 Introduction
The transition from lexical matching (e.g., BM25 Robertson and Zaragoza (2009)) to dense neural retrieval has revolutionized information discovery Karpukhin et al. (2020). However, deploying bi-encoder architectures in multi-tenant enterprise environments presents a "double scarcity" challenge. First, the Data Scarcity Bottleneck: enterprise tenants possess "dark data," proprietary corpora where relevance labels are non-existent, and standard benchmarks like BEIR Thakur et al. (2021) fail to capture the noisy, heterogeneous nature of these domains. Second, the Adaptation Latency Bottleneck: symmetric fine-tuning of both encoders Qu et al. (2021) incurs a massive "Re-indexing Tax," as any document encoder update necessitates re-generating embeddings for the entire corpus, computationally prohibitive for platforms hosting thousands of tenants.
@@
    DevRev Search Benchmark: A scalable pipeline that synthesizes training data without human annotators by pooling candidates from diverse retrievers via Reciprocal Rank Fusion Cormack et al. (2009) and employing LLM-as-a-Judge filtering Dai et al. (2023); Rahmani et al. (2024).
@@
    Zero-Reindexing Adaptation: An asymmetric fine-tuning strategy adapting only the query encoder via LoRA Hu et al. (2021), enabling tenant-specific adapters on a shared, frozen document index.
@@
    Layer Sensitivity Analysis: Empirical evidence that targeting specific transformer layers in the query encoder maximizes Recall while minimizing trainable parameters.
We validate our approach on DevRev Search and SciFact Wadden et al. (2022), demonstrating robust domain adaptation with a fraction of the cost of full fine-tuning.
@@ 2 Related Work
Synthetic Data.
Addressing label scarcity, recent works leverage LLMs for synthetic data generation Dai et al. (2023); Bonifacio et al. (2022); Wang et al. (2022). A key component is consistency filtering, discarding queries that fail to retrieve their source. We extend this by employing fusion-based candidate generation Cormack et al. (2009), aggregating diverse retrievers to minimize single-model bias and improve coverage.
LLM-as-Judge.
While LLMs show promise as relevance assessors Rahmani et al. (2024), concerns regarding bias persist Soboroff (2025). We mitigate this by using LLMs primarily for filtering pooled candidates, identifying positives rather than generating them, and validating a subset with human annotators.
Index-Preserving Adaptation.
Standard symmetric fine-tuning Karpukhin et al. (2020) incurs a prohibitive "Re-indexing Tax". Prior index-preserving approaches focus on pseudo-relevance feedback Yu et al. (2021) or asymmetric tuning Wang and Lyu (2023). We formalize this as Query-Side Adaptation, adapting the query manifold while freezing the document index.
Parameter-Efficient Fine-Tuning (PEFT).
PEFT methods like Adapters Houlsby et al. (2019) and LoRA Hu et al. (2021) have proven effective in retrieval Litschko et al. (2022). Unlike full fine-tuning, LoRA injects trainable low-rank matrices while freezing the backbone. Building on findings regarding intrinsic dimensionality Aghajanyan et al. (2020), we demonstrate that applying LoRA selectively to top query encoder layers maximizes efficiency for multi-tenant serving.
@@ 3 Dataset Generation
The recent paradigm shift in retrieval performance is largely attributed to the availability of high-quality, domain-specific data. However, a significant gap remains in the availability of publicly accessible enterprise search datasets that reflect the complex, semi-structured nature of real-world organizational data. Existing benchmarks like MS-MARCO or SciFact focus on web-scale passages or scientific abstracts, leaving a void for tasks involving technical support tickets, issue trackers, and internal documentation. By releasing the DevRev Search dataset, we aim to bridge this gap, providing the community with a high-fidelity benchmark for enterprise-specific retrieval.
Traditional manual annotation is not only prohibitively expensive and tedious but also suffers from low recall; human annotators cannot feasibly parse millions of documents, often leading to false negatives where relevant documents are overlooked simply because they were not reviewed. To address these challenges, we propose a scalable, automated pipeline to construct the DevRev Search dataset. Our methodology leverages a multi-stage process designed to maximize candidate coverage while maintaining high precision through the use of an LLM-as-judge.
@@ 3.1 Query Collection and Cleaning
We collected customer queries from production agent interactions as our source of real-world question data. However, raw customer queries often contain noise, including test queries, code snippets, and malformed inputs that are not legitimate natural language questions.
To ensure dataset quality, we implemented a multi-stage filtering process: (1) Length filtering: Removing queries with word counts in the bottom and top 25% percentiles. (2) Language detection: Retaining only English queries. (3) Deduplication: Removing exact duplicate queries. (4) Clustering-based diversity: Selecting representative samples from clusters to ensure diversity and avoid semantic repetition.
@@ 3.2 Multi-Retriever Annotation
Retrieval Ensemble: We applied an ensemble of seven diverse models: six dense retrievers (gemini-embedding-001 Google (2025), text-embedding-3-large OpenAI (2024), embed-english-v3 Cohere (2023), Qwen-3-Embedding-8B Zhang et al. (2025), GTE-Qwen2-7B-Instruct Li et al. (2023), SFR-Embedding-Mistral Meng et al. (2024)) and one lexical retriever (BM25), each returning the top 60 document chunks.
Union-based Aggregation: We computed the union of results from all retrievers to create a comprehensive candidate set of potentially relevant chunks. This union-based approach ensures that chunks retrieved by any of the seven models are included in the candidate set, maximizing coverage and recall across different retrieval paradigms yielding $\geq 60$ and $\leq 420$ unique candidate chunks per query across all models.
LLM-based Filtering: While the ensemble approach ensures high recall, it also introduces noise through false positives. To improve precision, we applied LLM-based filtering to the fused candidate set. Using a carefully designed prompt (provided in Appendix A.5), we tasked a large language model with identifying and retaining only the document chunks that genuinely contain information relevant to answering each query. This filtering step removes chunks that may have high lexical or semantic similarity to the query but lack substantive answer content.
@@
See Appendix A for further details on document segmentation and dataset statistics.
4 Experiments and Results
Standard bi-encoder optimization updates both query ($E_{q}$) and document ($E_{d}$) encoders. However, modifying $E_{d}$ incurs a re-indexing tax - the need to re-embed millions of documents and rebuild high-dimensional vector indices (e.g., HNSW). In large-scale, multi-tenant systems, this compute-intensive process introduces significant downtime and synchronization latency. To maintain a high velocity of model improvement without the prohibitive cost of index reconstruction, we propose Query-Side Adaptation: freezing the document encoder and index to enable near-instantaneous deployment.
We evaluate this strategy on two contrasting domains. First, our DevRev Search dataset (enterprise) features high relevance density (avg. 13.6 relevant chunks/query), testing the model’s capacity for broad semantic coverage and high recall. Conversely, SciFact Wadden et al. (2022) (scientific) exhibits low density (avg. 1.1 relevant docs/query), requiring identification of unique, specific evidence with high precision. We employ snowflake-arctic-embed-l-v2.0 Yu et al. (2024) and Qwen3-Embedding-4B Zhang et al. (2025) backbones. Training optimizes InfoNCE loss with 8 mined hard negatives per query essential for distinguishing subtly different technical concepts using a cosine learning rate scheduler. Refer to A.6 for our detailed experimentation setup.
Our analysis covers: (1) Query-Only vs. Joint Tuning to quantify the performance trade-off of freezing the index; (2) LoRA Approximation and Scaling to analyze sensitivity to rank $r$; and (3) Module Targeting to identify which transformer components (e.g., Attention vs. MLP) yield the highest returns.
4.1 Comparison of Fine-Tuning Strategies
We first address whether query-encoder-only fine-tuning ($Q$) can achieve performance parity with joint Query-Document fine-tuning ($QD$). The results obtained by using the optimal hyperparameter settings for each configuration, are summarized in Fig 1.
DevRev Search Results.
On the DevRev Search benchmark surprisingly, Query-Only ($Q$) consistently outperforms Query-Document ($QD$) on the enterprise benchmark. We attribute this to asymmetric regularization. In low-resource specialized domains, joint optimization effectively doubles the parameter space, increasing the risk of overfitting and distorting the pre-trained document manifold. By freezing the document encoder, we enforce a structural constraint projecting queries into a stable target space which improves generalization to unseen test queries.
SciFact Results.
On the SciFact dataset, $QD$ remains the upper bound, confirming that complex scientific alignment benefits from reshaping both spaces. However, $Q$ remains highly competitive, recovering the vast majority of performance gains (within 1-2% of $QD$). This establishes Query-Only adaptation as a Pareto-optimal strategy for production: it delivers comparable accuracy to joint tuning while completely eliminating the prohibitive re-indexing tax.
(a) DevRev Search
(b) SciFact
Figure 1: Comparison of Recall@10 for Baseline, Query-Document (QDQD), and Query-Only ($Q$) fine-tuning. On DevRev Search, $Q$ surprisingly outperforms QDQD, while on SciFact, $Q$ remains highly competitive.
(a) DevRev Search
(b) SciFact
Figure 2: Performance vs. LoRA Rank. Note Qwen’s preference for lower ranks ($r=8$) on DevRev Search versus mid-ranks ($r=64$) on SciFact.
4.2 Impact of LoRA Rank
We next investigate the sensitivity of the models to the LoRA rank ($r$), which controls the capacity of the trainable adapters. As shown in Fig 2, the optimal rank for DevRev Search varies significantly by model architecture. In particular, since the DevRev Search dataset is relatively small, larger models like Qwen3-Embedding-4B tend to overfit on higher ranks. On SciFact however, we observe that higher ranks are generally preferred.
4.3 Targeted Lora Module Fine-tuning
Finally, we analyze which transformer sub-layers (Query-Value QV, Feed-Forward Network FFN, Query-Key-Value QKV, or All Layers) yield the best adaptation results.
(a) DevRev Search
(b) SciFact
Figure 3: Targeted Module Fine Tuning
The findings from Fig 3 highlight that the optimal fine-tuning strategy is highly dependent on the model-dataset pair. While the smaller Snowflake model benefits from maximizing capacity (Higher Rank, All Layers), the larger Qwen model performs well with targeted regularization (Lower Rank, QV/FFN modules) to achieve peak retrieval quality, indicating a clear trend.

### v2

# Succeeding at Scale: Automated Dataset Construction and Query-Side Adaptation for Multi-Tenant Search

@@ Abstract
Large-scale multi-tenant retrieval systems generate extensive query logs but lack curated relevance labels for effective domain adaptation, resulting in substantial underutilized “dark data”. This challenge is compounded by the high cost of model updates, as jointly fine-tuning query and document encoders requires full corpus re-indexing, which is impractical in multi-tenant settings with thousands of isolated indices. We introduce DevRev-Search [^2] https://huggingface.co/datasets/devrev/search, a passage retrieval benchmark for technical customer support built via a fully automated pipeline. Candidate generation uses fusion across diverse sparse and dense retrievers, followed by an LLM-as-a-Judge for consistency filtering and relevance labeling. We further propose an Index-Preserving Adaptation strategy that fine-tunes only the query encoder, achieving strong performance gains while keeping document indices fixed. Experiments on DevRev-Search, SciFact, and FiQA-2018 show that Parameter-Efficient Fine-Tuning (PEFT) of the query encoder delivers a remarkable quality–efficiency trade-off, enabling scalable and practical enterprise search adaptation.
@@ 1 Introduction
The transition from lexical matching methods such as BM25 (Robertson and Zaragoza (2009)) to dense neural retrieval has revolutionized information discovery (Karpukhin et al. (2020)). However, deploying bi-encoder architectures in multi-tenant enterprise environments presents a ”double scarcity” challenge. First, the Data Scarcity Bottleneck: enterprise tenants possess ”dark data,” proprietary corpora where relevance labels are non-existent, and standard benchmarks like BEIR (Thakur et al. (2021)) fail to capture the noisy, heterogeneous nature of these domains. Second, the Adaptation Latency Bottleneck: symmetric fine-tuning of both encoders (Qu et al. (2021)) incurs a massive ”Re-indexing Tax,” as any document encoder update necessitates re-generating embeddings for the entire corpus, computationally prohibitive for platforms hosting thousands of tenants.
@@
    DevRev-Search Benchmark: A scalable pipeline that synthesizes training data without human annotators by pooling candidates from diverse retrievers and employing LLM-as-a-Judge filtering (Dai et al. (2023); Rahmani et al. (2024)).
@@
    Zero-Reindexing Adaptation: An asymmetric fine-tuning approach that adapts only the query encoder, enabling tenant-specific models with a frozen document index.
@@
    Parameter Efficient Query Adaptation: We show that parameter-efficient techniques such as LoRA (Hu et al. (2022)), embedding transformation using linear and feed-forward projections, and fine-tuning of only limited transformer layers—can achieve performance that matches or closely approaches full query adaptation. We further conduct ablations over LoRA ranks and modules, as well as the number of unfrozen transformer layers.
We validate our approach on DevRev-Search, SciFact (Wadden et al. (2022)), and FiQA-2018 (Maia et al. (2018)), demonstrating robust domain adaptation with a fraction of the cost of full bi-encoder fine-tuning.
@@ 2 Related Work
Synthetic Data: Addressing label scarcity, recent works leverage LLMs for synthetic data generation (Dai et al. (2023); Bonifacio et al. (2022); Wang et al. (2022)). A key component is consistency filtering, discarding queries that fail to retrieve their source. We extend this by employing fusion-based candidate generation (Cormack et al. (2009)), aggregating diverse retrievers to minimize single-model bias and improve coverage. LLM-as-Judge: While LLMs show promise as relevance assessors (Rahmani et al. (2024)), concerns regarding bias persist (Soboroff (2025)). We mitigate this by using LLMs primarily for filtering pooled candidates, identifying positives rather than generating them, and validating a subset with human annotators.
Index-Preserving Adaptation: Standard symmetric fine-tuning (Karpukhin et al. (2020)) incurs a prohibitive ”Re-indexing Tax”. Prior index-preserving approaches focus on pseudo-relevance feedback (Yu et al. (2021)) or asymmetric tuning (Wang and Lyu (2023)). We formalize this as Query-Only Adaptation, adapting the query manifold while freezing the document index. Parameter-Efficient Fine-Tuning (PEFT): PEFT methods like Adapters (Houlsby et al. (2019)), LoRA (Hu et al. (2022)), embedding projections (Yoon et al. (2024)) and partial unfreezing of layers (Lee et al. (2022)) have proven effective in retrieval (Litschko et al. (2022)). Unlike full fine-tuning, PEFT methods use much fewer trainable parameters compared to full fine-tuning. Building on findings regarding intrinsic dimensionality (Aghajanyan et al. (2020)), we demonstrate that applying PEFT to the query encoder results in both high quality and efficiency for multi-tenant serving.
@@ 3 Dataset Generation
Recent retrieval gains are driven by high-quality, domain-specific data, yet publicly available enterprise search datasets remain scarce. Existing benchmarks such as MS-MARCO and SciFact focus on web passages or scientific abstracts and do not capture the semi-structured, heterogeneous nature of enterprise data, including support tickets, issue trackers, and internal documentation. DevRev-Search addresses this gap by providing a high-fidelity benchmark for enterprise-specific retrieval.
Manual annotation is costly, slow, and inherently low-recall, as annotators cannot exhaustively review large corpora, resulting in systematic false negatives. To address this, we introduce a scalable, automated dataset construction pipeline that maximizes candidate coverage through multi-stage retrieval while maintaining high precision via an LLM-as-judge.
@@ 3.1 Query Collection and Cleaning
We collected customer queries from production agent interactions as our source of real-world question data. However, raw customer queries often contain noise, including test queries, code snippets, and malformed inputs that are not legitimate natural language questions. To ensure dataset quality, we implemented a multi-stage filtering process: (1) Length filtering: Removing queries with word counts in the bottom and top 25% percentiles. (2) Language detection: Retaining only English queries. (3) Deduplication: Removing exact duplicate queries. (4) Clustering-based diversity: Selecting representative samples from clusters to ensure diversity and avoid semantic repetition.
3.2 Document Segmentation and Semantic Granularity
Enterprise documents are typically long and sparsely relevant to specific queries, which poses challenges for dense retrieval due to encoder token limits and the limited expressive capacity of fixed-size embeddings. Encoding entire documents often results in diluted representations that obscure fine-grained relevance. To address this, we collected all public documentation of DevRev and applied Recursive Character Splitting LangChain (2025). Documents are segmented into chunks of up to 500 characters with no overlap, maximizing semantic coverage while preserving precision. The recursive strategy favors natural structural boundaries (e.g., paragraphs, sentences, then spaces), ensuring each fragment is both embedding-friendly and semantically self-contained.
@@ 3.3 Multi-Retriever Annotation
Retrieval Ensemble: We applied an ensemble of seven diverse models: six dense retrievers (gemini-embedding-001 Google (2025), text-embedding-3-large (OpenAI (2024)), embed-english-v3 (Cohere (2023)), Qwen-3-Embedding-8B (Zhang et al. (2025)), GTE-Qwen2-7B-Instruct (Li et al. (2023)), SFR-Embedding-Mistral (Meng et al. (2024))) and one lexical retriever (BM25) ( Robertson and Zaragoza (2009)), each returning the top 60 document chunks.
Union-based Aggregation: We computed the union of results from all retrievers to create a comprehensive candidate set of potentially relevant chunks. This ensures that chunks retrieved by any of the seven models are included in the candidate set, maximizing coverage and recall across different retrieval paradigms yielding $\geq 60$ and $\leq 420$ unique candidate chunks per query across all models.
LLM-based Filtering: While the ensemble approach boosts recall, it introduces noise via false positives. To improve precision, we applied LLM-based filtering to the fused candidate set. Using a carefully designed prompt (Appendix A.3), a large language model was tasked with retaining only document chunks genuinely relevant to each query, removing those with high lexical or semantic similarity but lacking substantive answer content.
@@
3.3.1 Analysis of Retriever Contributions
To assess the necessity of a multi-retriever ensemble, we analyzed the individual performance, overlap, and complementarity of each retriever.
Individual Recall.
We measured standalone recall by retrieving the top 420 candidates per query from each model and comparing them to the final ground truth. As shown in Table 3 (Appendix A.2), even the best-performing retriever (gemini-embedding-001) achieves only 82.48 recall, showing that any single retriever misses a substantial portion of relevant documents and introduces bias.
Redundancy and Diversity.
To evaluate redundancy, we performed a leave-one-out ablation: for each retriever $M_{i}$, we aggregated candidates from the remaining six models and measured recall against the ground truth. If $\text{All}\setminus{M_{i}}$ reached perfect recall, $M_{i}$ would be redundant. However, as reported in Table 4 (Appendix A.2), recall ranges from 93.25 to 97.13, demonstrating that every retriever contributes unique relevant candidates. While removing strong dense models causes the largest drop, excluding BM25 also reduces coverage, highlighting the value of a hybrid ensemble.
Overall, our pipeline balances recall-oriented aggregation with precision-oriented filtering. The seven-retriever ensemble provides broad coverage by combining dense semantic and lexical signals, while LLM-based filtering applies deep reasoning to eliminate false positives and finalize high-quality labels. See Appendix A.1 for additional dataset statistics.
4 Query-Only Adaptation
Standard bi-encoder fine-tuning updates both the query ($E_{q}$) and document ($E_{d}$) encoders, but modifying $E_{d}$ requires costly re-embedding and index reconstruction (e.g., HNSW) over millions of documents, which is impractical in large-scale, multi-tenant systems. To eliminate this re-indexing cost while enabling rapid iteration, we propose Query-Only Adaptation, which freezes the document encoder and index for very fast deployment.
We evaluate this approach across three datasets with varying relevance densities: DevRev-Search (enterprise, high density; 13.6 relevant chunks/query), SciFact (scientific, low density; 1.1), and FiQA-2018 (financial, medium density; 2.6). These respectively emphasize recall, precise evidence retrieval, and balanced precision–recall. We use snowflake-arctic-embed-l-v2.0 Yu et al. (2024) and Qwen3-Embedding-4B Zhang et al. (2025) as backbone encoders.
Training employs InfoNCE loss with mined hard negatives to improve discrimination among semantically similar documents. A single tuned temperature is used throughout training, and stability is ensured via asynchronous ANCE training (Xiong et al., 2020) to prevent representation collapse. Additional experimental details are provided in Appendix A.4.
Our analysis focuses on: (1) Query-Only vs. Query–Document Fine-Tuning, measuring the performance impact of freezing the document index; and (2) Parameter-Efficient Query Adaptation, evaluating whether low-rank and parameter-efficient methods can match full fine-tuning while enabling cost-effective multi-tenant deployment.
4.1 Query Only vs Query-Document Fine-Tuning
We first evaluate whether asymmetric Query-Only fine-tuning ($Q$) can reach performance parity with symmetric Query-Document fine-tuning ($QD$). Table 1 summarizes the results across both architectures. The data demonstrates that freezing the document encoder incurs minimal performance loss. In all cases, the $Q$ strategy significantly improves over the untrained baseline. Crucially, it remains highly competitive with the computationally expensive $QD$ approach, achieving comparable retrieval metrics across all domains and, in the case of qwen3-4b on SciFact, even marginally outperforming the joint tuning strategy in Recall@10.
Table 1: Comparison of Recall@10 and NDCG@10 in retrieval performance for baseline untrained model (Base) with Query-Document (QD), and Query-Only (Q) finetuning
Model Variant DevRev-Search SciFact FiQA
recall@10 ndcg@10 recall@10 ndcg@10 recall@10 ndcg@10
arctic-l-v2 Base 0.256 0.304 0.819 0.691 0.533 0.459
QD 0.314 0.362 0.869 0.748 0.578 0.505
Q 0.296 0.343 0.854 0.727 0.554 0.478
qwen3-4b Base 0.219 0.264 0.903 0.769 0.649 0.571
QD 0.325 0.396 0.949 0.875 0.712 0.628
Q 0.327 0.370 0.953 0.830 0.688 0.602
4.2 Parameter Efficient Fine-Tuning
We examine whether Query-Only full fine-tuning performance can be achieved using PEFT methods, reducing trainable parameters and enabling scalable multi-tenant adaptation. We evaluate several PEFT techniques: LoRA, a linear projection head, a 2 hidden-layer feed-forward network (FFN) on embeddings, and unfreezing the top 8 transformer layers. The number and proportion of trainable parameters for each method are reported in Table 5 (Appendix A.4).
Results in Table 2 show that the best LoRA configuration consistently matches or outperforms full fine-tuning, likely due to its implicit regularization via layer-wise residual learning. Linear and FFN heads are also competitive; notably, FFN surpasses full fine-tuning on SciFact. In contrast, fine-tuning only the top 8 transformer layers underperforms despite using more parameters, highlighting LoRA’s effectiveness as a PEFT strategy.
Table 2: Comparison of retrieval performance for various parameter-efficient methods with full finetuning for query-only adaptation (Here *Lora and 8-Tr refer to the best LoRA configuration and fine-tuning only 8 transformer layers at the top of the base model, respectively)
Model Type DevRev-Search SciFact FiQA
recall@10 ndcg@10 recall@10 ndcg@10 recall@10 ndcg@10
arctic-l-v2 Linear 0.271 0.314 0.849 0.722 0.543 0.470
FFN 0.281 0.318 0.880 0.743 0.534 0.464
*LoRA 0.309 0.342 0.915 0.827 0.555 0.478
8-Tr 0.273 0.317 0.849 0.718 0.547 0.476
Full 0.296 0.343 0.854 0.727 0.554 0.478
qwen3-4b Linear 0.331 0.358 0.935 0.815 0.666 0.583
FFN 0.340 0.362 0.946 0.861 0.660 0.577
*LoRA 0.355 0.399 0.950 0.844 0.694 0.607
8-Tr 0.312 0.356 0.932 0.802 0.673 0.590
Full 0.327 0.370 0.953 0.830 0.688 0.602
Figure 1: Ablation results for query only adapation for DevRev-Search across (a) varying LoRA ranks (with $\alpha=r$), (b) different combinations of target LoRA modules, and (c) varying number of top unfrozen transformer layers with the rest of the model frozen
LoRA Rank Ablations.
We study sensitivity to the LoRA rank ($r$) with $\alpha=r$. On DevRev-Search (Fig. 1(a)), arctic-l-v2 improves monotonically and saturates at high ranks, whereas qwen3-4b shows a peaked profile with best performance at $r=32$, likely due to overfitting at larger ranks. Results on SciFact and FiQA (Table 7, Appendix A.5) show a consistent pattern: $r\in[32,64]$ provides the best trade-off between capacity and regularization. Although $r=128$ can sometimes yield minor gains, performance largely saturates, making $32$ or $64$ the most efficient choices.
Lora Target Module Ablations.
We also compare LoRA-based query-only adaptation across transformer sub-layer combinations — dense, QV (Query-Value), QKV (Query-Key-Value), and all modules. Figure 1(b) shows the DevRev-Search results. For results on the other datasets, please refer to Table 8 (Appendix A.5). While generally taking all layers is the best choice, we observe that just tuning dense layers provides an impressive performance-efficiency tradeoff compared to other modules. This observation is more visible with qwen3-4b.
Transformer Layer Ablations.
Finally, we freeze the base model and ablate the number of transformer layers at the top that are unfrozen. As seen in Fig 1(c), the performance monotonically increases with the number of unfrozen transformer layers, regardless of model architecture, on DevRev-Search. Table 9 (Appendix A.5) reports the results for the other datasets, which exhibit the same trend.